# ./references.bib

### v1

@@ Abstract
This paper introduces Gauge Flow Models, a novel class of Generative Flow Models. These models incorporate a learnable Gauge Field within the Flow Ordinary Differential Equation (ODE). A comprehensive mathematical framework for these models, detailing their construction and properties, is provided. Experiments using Flow Matching on Gaussian Mixture Models demonstrate that Gauge Flow Models yields significantly better performance than traditional Flow Models of comparable or even larger size. Additionally, unpublished research indicates a potential for enhanced performance across a broader range of generative tasks.
@@ A Gauge Flow Model is defined within the mathematical framework of associated bundles:
where $P=(P,G,M,\pi_{P})$ is a principal bundle with structure group $G$ and $F$ is a $G$-space. The dynamics are governed by the following neural ODE:
^$\hat{\nabla}_{dt}x(t):=v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu\nu}(x(t),t)d^{\mu}(x(t),t)v^{\nu}(x(t),t))$
@@
  • $A_{\mu\nu}(x(t),t)$ is a learnable Gauge Field valued in the Lie algebra of the Gauge Group $G$.
@@
  • $v^{\nu}(x(t),t))$ is a further learnable vector field corresponding to a section on the associated bundle $P\times_{G}F$.
@@
The novelty of this Flow ODE lies in the inclusion of the Gauge Term, which introduces a novel combination of learnable and non-learnable vector fields – a feature absent from standard Neural Flow Models. The standard Neural Flow Model, as defined by [CNFs, FM], is characterized by:
@@ 2   Mathematical Background
  • Differential Geometry: [DifferentialGeometry, DifferentialGeometryManifoldConnections, DifferentialGeometryManifold, SmoothManifolds, Manifolds]
  • Gauge Theory: [MathematicalGaugeTheory, GeometryTopologyPhysics, FibreBundle]
  • Lie Groups and Lie Algebras: [LieGroupsLieAlgebras, LieAlgebras]
@@ 2.1   Fiber Bundle
[…]
@@ 2.1.2   Principal Bundle
:$\bullet$:∙ : $G\times G\rightarrow G$
@@
Lie Algebra A Lie algebra $\mathfrak{g}$ of a Lie Group $G$ is simply the tangent space $T_{e}G$ at the identity element $e$ to the Lie Group. This vector space $\mathfrak{g}$ inherits a bilinear bracket [,][-,-][ - , - ] from the group structure of $G$:
$[-,-]:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$
@@
Connection on a Principal Bundle A connection on a principal bundle $P=(P,B,G,\pi)$ is a connection $HP\subset TP$ on the fiber bundle $P$ with characterized by the G-equivariance condition:
@@ 2.1.3   Associated Bundle
with principal bundle $P=(P,G,B,\pi)$ and $F$ being a smooth manifold (or vector space) equipped with a smooth left action of Lie group $G$.
@@
${P\times F}$${\hat{A}=P\times_{G}F}$${P}$${B}$$\scriptstyle{q}$$\scriptstyle{\mathrm{pr}_{P}}$$\scriptstyle{\pi_{\hat{A}}}$$\scriptstyle{\pi}$
@@ A Gauge Flow Model is defined within the framework of the associated bundle $\hat{A}$:
  • The principal bundle $P=(P,G,M,\pi_{P})$ with Lie group $G$ and base manifold $M$.
@@
^$\hat{\nabla}_{dt}x(t):=v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu\nu}(x(t),t)d^{\mu}(x(t),t)v^{\nu}(x(t),t))$
@@
  • $A$ is the learnable gauge field corresponding to the connection on the principal bundle $P$, also represented by a Neural Network.
@@
  • $v^{\nu}(x(t),t):M\rightarrow P\times_{G}F$ is the fiber section, which can be learned by a Neural Network.
@@
Training a Gauge Flow Model employs the Riemannian Flow Matching (RFM) framework [RFM], which generalizes ordinary Flow Matching (FM) [FM]. The training objective is the Gauge Flow Model loss, defined as:
$\mathcal{L}_{\mathrm{GFM}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ x\sim p_{t}\end{subarray}}\bigg{\lVert}\Big{[}v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu\nu}(x(t),t)d^{\mu}(x(t),t)v^{\nu}(x(t),t))\Big{]}-u_{t}(x)\bigg{\rVert}_{g_{x}}^{2}$
@@
  • Probability-density path $p_{t}:M\to\mathbb{R}_{+}$, $t\in[0,1]$, as defined in [RFM]. For each $t$, $p_{t}$ satisfies
[…]
@@
[…]
@@
Computing the Gauge Flow Model (GFM) loss $\mathcal{L}_{\mathrm{GFM}}$ exactly is generally intractable. However, this loss is equivalent to the Riemannian Flow Matching (RFM) objective. To address this intractability, [RFM] introduces the Riemannian Conditional Flow Matching (RCFM) loss—a single-sample, unbiased Monte Carlo estimator of the RFM objective. The key idea is to replace the marginal target field $u_{t}(x)$ with its conditional counterpart $u_{t}(x\mid x_{1})$, enabling tractable training. This formulation allows for efficient, simulation-free learning on simple geometries and scalable neural network training on general Riemannian manifolods. Notably, this approach can also be applied to train Gauge Flow Models.
For full derivations and implementation details, refer to [RFM, FM].
@@ 4.1   Model
  • The trivial principal bundle $P=(P=M\times G,G,M,\pi)$ with structure Lie group $G=SO(N)$ and base manifold $M=\mathbb{R}^{N}$.
@@
  • The directional vector field $d^{\mu}(x(t),t)$ is either identified with the vector field $v_{\theta}(x(t),t)$ or $v^{\nu}(x(t),t)$ leading to two distinct Gauge Flow Model variants.
  • The vector field $v^{\nu}(x(t),t):M\rightarrow P\times_{G}F$ is expressed in the gauge defined by orthonormal bases of $\mathbb{R}^{N}$. In this gauge, the field $v^{\nu}(x(t),t)$ corresponds to a global map $M\rightarrow F$, which is modeled using a Neural Network.
@@
The layer width depends on the dimension $N$ and is set to 32 for $N>10$ and to 64 otherwise. Here, SiLU refers to the Sigmoid Linear Unit activation function. The Plain Flow Model is implemented as a multilayer perceptron (MLP), where the number of layers and their widths are functions of the input dimension $N$. Initially, the network architecture follows the layer dimensions $[N+1,128,128,128,N]$, with SiLU used as the activation function throughout. These Gauge Flow Models are generally trained within the framework of Riemannian Flow Matching [RFM]. However, since the base manifold is $\mathbb{R}^{N}$, training can be performed using the standard Flow Matching [FM].
@@ 4.2   Dataset
[…]
@@ 4.3.1   Train Loss
The Gauge Flow Model utilizing the directional vector field $v_{\theta}(x(t),t)$ or $v^{\nu}(x(t),t)$, demonstrated superior training performance. To facilitate meaningful comparisons, the training loss was normalized relative to the loss of the $v_{\theta}(x(t),t)$ Gauge Flow Models. Notably, the training performance varied across different Lie group dimensions ($N$s), with the Gauge Flow Model using $v_{\theta}(x(t),t)$ consistently exhibiting the best results for most values of $N$.
Figure 3: Train Loss Comparison (lower is better): The normalized train loss is plotted against the dimension $N$ to compare the performance of the Gauge Flow Model using directional vector fields $v_{\theta}(x(t),t)$ and $v^{\nu}(x(t),t)$ with the standard plain Flow Model.
@@ 4.3.2   Test Loss
Figure 4: Test Loss Comparison (lower is better): The normalized test loss is plotted against the dimension $N$ to compare the performance of the Gauge Flow Model using directional vector fields $v_{\theta}(x(t),t)$ and $v^{\nu}(x(t),t)$ with the standard plain Flow Model.
@@ 4.3.3   Number of Parameters
Figure 5: Number of Parameters: A comparison of the number of parameters for the plain Flow Model and the Gauge Flow Models across various dimensions $N$.
@@ 5   Related Work
The concept of Continuous Normalizing Flows (CNFs), as introduced in [CNFs], traditionally relies on computationally intensive Ordinary Differential Equation (ODE) solvers during training, guided by a maximum likelihood objective. While effective, this training procedure scales poorly due to the repeated integration of ODEs. To mitigate this limitation, the Flow Matching paradigm was proposed in [FM]. Instead of maximizing likelihood, Flow Matching formulates training as an $L_{2}$ regression problem that directly aligns a learnable vector field with a target velocity field. This target velocity is derived from known boundary conditions, significantly reducing computational overhead during training. Initially formulated in Euclidean space, Flow Matching was later extended to Riemannian manifolds in the Riemannian Flow Matching (RFM) framework [RFM]. In RFM, the vector field matching operates on curved spaces while respecting the geometry induced by a fixed Riemannian metric. However, it is important to note that the connections—i.e., the geometric structures defining parallel transport and differentiation—are determined by the manifold and are not themselves learnable parameters.
Parallel to these developments, several lines of research have explored the connections between gauge theory and deep learning. A notable example is presented in [GNN], where a gauge-invariant framework is introduced that encodes invariance directly into the inputs by constructing gauge-invariant features before passing them through standard neural architectures. This approach imposes gauge symmetry directly on the inputs instead of embedding it in the model dynamics.
Another approach leveraging geometric and gauge-theoretic principles is presented in [GCNNs], where standard Convolutional Neural Networks (CNNs) are generalized to operate on arbitrary manifolds by enforcing coordinate independence. This work builds upon the foundations laid by steerable CNNs [SteerableCNNs] and Group Equivariant CNNs (G-CNNs) [GECNNs], extending the principle of equivariance to manifold-valued data. Further advancing this direction, Lattice Gauge Equivariant Convolutional Neural Networks (L-CNNs) [LGCNN] incorporate local gauge symmetry directly into the network architecture. These models integrate gauge equivariance into both convolutional layers and bilinear operations used to compute Wilson loops, which are gauge-invariant observables from lattice gauge theory. This framework was subsequently extended in [GLCNN] to include global lattice symmetries, such as rotations and reflections, thereby achieving a comprehensive symmetry-aware neural network design.
[…]

### v2

# Gauge Flow Models

@@ Abstract
This paper introduces Gauge Flow Models, a novel class of Generative Flow Models. These models incorporate a learnable Gauge Field within the Flow Ordinary Differential Equation (ODE). A comprehensive mathematical framework for these models, detailing their construction and properties, is provided. Experiments using Flow Matching on Gaussian Mixture Models demonstrate that Gauge Flow Models yield significantly better performance than traditional Flow Models of comparable or even larger size. Additionally, unpublished research indicates a potential for enhanced performance across a broader range of generative tasks.
@@ A Gauge Flow Model is defined within the mathematical framework of associated bundles:
where $P=(P,M,G,\pi_{P})$ is a principal bundle with structure group $G$ and $F$ is a $G$-space. The dynamics are governed by the following neural ODE:
$\hat{\nabla}_{dt}x(t):=v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))$
@@
  • $A_{\mu}(x(t),t)$ is a learnable Gauge Field valued in the Lie algebra of the Gauge Group $G$.
@@
  • $v_{s}(x(t),t)$ is a further learnable vector field corresponding to a section on the associated bundle $P\times_{G}F$.
@@
The novelty of this Flow ODE lies in the inclusion of the Gauge Term, which introduces a novel combination of learnable and non-learnable vector fields – a feature absent from standard Neural Flow Models. The standard Neural Flow Model, as defined by [1, 2], is characterized by:
@@ 2  Mathematical Background
  • Differential Geometry: [3, 4, 5, 6, 7]
  • Gauge Theory: [8, 9, 10]
  • Lie Groups and Lie Algebras: [11, 12]
@@ 2.1  Fiber Bundle
$\pi^{-1}(U_{x})$$U\times F$$U_{x}$$U_{x}$$\scriptstyle{\Psi}$$\scriptstyle{\pi}$$\scriptstyle{\mathrm{pr}_{1}}$$\scriptstyle{\mathrm{id}_{U_{x}}}$
@@ 2.1.2  Principal Bundle
$\bullet:$ $G\times G\rightarrow G$
@@
Lie Algebra A Lie algebra $\mathfrak{g}$ of a Lie Group $G$ is simply the tangent space $T_{e}G$ at the identity element $e$ to the Lie Group. This vector space $\mathfrak{g}$ inherits a bilinear bracket [,][-,-] from the group structure of $G$:
$\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$
@@
Connection on a Principal Bundle A connection on a principal bundle $P=(P,B,G,\pi)$ is a connection $HP\subset TP$ on the fiber bundle $P$ characterized by the G-equivariance condition:
@@ 2.1.3  Associated Bundle
with principal bundle $P=(P,B,G,\pi)$ and $F$ being a smooth manifold (or vector space) equipped with a smooth left action of Lie group $G$.
@@
[…]
@@ A Gauge Flow Model is defined within the framework of the associated bundle $\hat{A}$:
  • The principal bundle $P=(P,M,G,\pi_{P})$ with Lie group $G$ and base manifold $M$.
@@
$\hat{\nabla}_{dt}x(t):=v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))$
@@
  • AA is the learnable gauge field corresponding to the connection on the principal bundle $P$, also represented by a Neural Network.
@@
  • $v_{s}(x(t),t):M\rightarrow P\times_{G}F$ is the fiber section, which can be learned by a Neural Network.
@@
Training a Gauge Flow Model employs the Riemannian Flow Matching (RFM) framework [13], which generalizes ordinary Flow Matching (FM) [2]. The training objective is the Gauge Flow Model loss, defined as:
$\mathcal{L}_{\mathrm{GFM}}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ x\sim p_{t}\end{subarray}}\bigg{\lVert}\Big{[}v_{\theta}(x(t),t)-\alpha(t)\Pi_{M}(A_{\mu}(x(t),t)d^{\mu}(x(t),t)v_{s}(x(t),t))\Big{]}-u_{t}(x)\bigg{\rVert}_{g_{x}}^{2}$
@@
  • Probability-density path $p_{t}:M\to\mathbb{R}_{+}$, $t\in[0,1]$, as defined in [13]. For each $t$, $p_{t}$ satisfies
    $\int_{M}p_{t}(x)\,d\mathrm{vol}_{x}=1,$
@@
    $u_{t}(x)=\int_{M}u_{t}\!\bigl{(}x\mid x_{1}\bigr{)}\,\frac{p_{t}\!\bigl{(}x\mid x_{1}\bigr{)}\,q(x_{1})}{p_{t}(x)}\,d\mathrm{vol}_{x_{1}},$
@@
Computing the Gauge Flow Model (GFM) loss $\mathcal{L}_{\mathrm{GFM}}$ exactly is generally intractable. However, this loss is equivalent to the Riemannian Flow Matching (RFM) objective. To address this intractability, [13] introduces the Riemannian Conditional Flow Matching (RCFM) loss—a single-sample, unbiased Monte Carlo estimator of the RFM objective. The key idea is to replace the marginal target field $u_{t}(x)$ with its conditional counterpart $u_{t}(x\mid x_{1})$, enabling tractable training. This formulation allows for efficient, simulation-free learning on simple geometries and scalable neural network training on general Riemannian manifolds. Notably, this approach can also be applied to train Gauge Flow Models.
For full derivations and implementation details, refer to [13, 2].
@@ 4.1  Model
  • The trivial principal bundle $P=(P=M\times G,M,G,\pi)$ with structure Lie group $G=SO(N)$ and base manifold $M=\mathbb{R}^{N}$.
@@
  • The directional vector field $d^{\mu}(x(t),t)$ is either identified with the vector field $v_{\theta}(x(t),t)$ or $v_{s}(x(t),t)$ leading to two distinct Gauge Flow Model variants.
  • The vector field $v_{s}(x(t),t):M\rightarrow P\times_{G}F$ is expressed in the gauge defined by orthonormal bases of $\mathbb{R}^{N}$. In this gauge, the field $v_{s}(x(t),t)$ corresponds to a global map $M\rightarrow F$, which is modeled using a Neural Network.
@@
The layer width depends on the dimension $N$ and is set to 32 for $N>10$ and to 64 otherwise. Here, SiLU refers to the Sigmoid Linear Unit activation function. The Plain Flow Model is implemented as a multilayer perceptron (MLP), where the number of layers and their widths are functions of the input dimension $N$. Initially, the network architecture follows the layer dimensions $[N+1,128,128,128,N]$, with SiLU used as the activation function throughout. These Gauge Flow Models are generally trained within the framework of Riemannian Flow Matching [13]. However, since the base manifold is $\mathbb{R}^{N}$, training can be performed using the standard Flow Matching [2].
@@ 4.2  Dataset
Sampling:  Draw $k\sim\mathrm{Cat}(\pi)$ and $x\sim\mathcal{N}\!\bigl{(}\mu_{k},\,0.5\,I_{N}\bigr{)}.$
@@ 4.3.1  Train Loss
The Gauge Flow Model utilizing the directional vector field $v_{\theta}(x(t),t)$ or $v_{s}(x(t),t)$, demonstrated superior training performance. To facilitate meaningful comparisons, the training loss was normalized relative to the loss of the $v_{\theta}(x(t),t)$ Gauge Flow Models. Notably, the training performance varied across different Lie group dimensions ($N$s), with the Gauge Flow Model using $v_{\theta}(x(t),t)$ consistently exhibiting the best results for most values of $N$.
Figure 3: Train Loss Comparison (lower is better): The normalized train loss is plotted against the dimension NN to compare the performance of the Gauge Flow Model using directional vector fields $v_{\theta}(x(t),t)$ and $v_{s}(x(t),t)$ with the standard plain Flow Model.
@@ 4.3.2  Test Loss
Figure 4: Test Loss Comparison (lower is better): The normalized test loss is plotted against the dimension NN to compare the performance of the Gauge Flow Model using directional vector fields $v_{\theta}(x(t),t)$ and $v_{s}(x(t),t)$ with the standard plain Flow Model.
@@ 4.3.3  Number of Parameters
Figure 5: Number of Parameters: A comparison of the number of parameters for the plain Flow Model and the Gauge Flow Models across various dimensions NN.
@@ 5  Related Work
The concept of Continuous Normalizing Flows (CNFs), as introduced in [1], traditionally relies on computationally intensive Ordinary Differential Equation (ODE) solvers during training, guided by a maximum likelihood objective. While effective, this training procedure scales poorly due to the repeated integration of ODEs. To mitigate this limitation, the Flow Matching paradigm was proposed in [2]. Instead of maximizing likelihood, Flow Matching formulates training as an $L_{2}$ regression problem that directly aligns a learnable vector field with a target velocity field. This target velocity is derived from known boundary conditions, significantly reducing computational overhead during training. Initially formulated in Euclidean space, Flow Matching was later extended to Riemannian manifolds in the Riemannian Flow Matching (RFM) framework [13]. In RFM, the vector field matching operates on curved spaces while respecting the geometry induced by a fixed Riemannian metric. However, it is important to note that the connections—i.e., the geometric structures defining parallel transport and differentiation—are determined by the manifold and are not themselves learnable parameters.
Parallel to these developments, several lines of research have explored the connections between gauge theory and deep learning. A notable example is presented in [14], where a gauge-invariant framework is introduced that encodes invariance directly into the inputs by constructing gauge-invariant features before passing them through standard neural architectures. This approach imposes gauge symmetry directly on the inputs instead of embedding it in the model dynamics.
Another approach leveraging geometric and gauge-theoretic principles is presented in [15], where standard Convolutional Neural Networks (CNNs) are generalized to operate on arbitrary manifolds by enforcing coordinate independence. This work builds upon the foundations laid by steerable CNNs [16] and Group Equivariant CNNs (G-CNNs) [17], extending the principle of equivariance to manifold-valued data. Further advancing this direction, Lattice Gauge Equivariant Convolutional Neural Networks (L-CNNs) [18] incorporate local gauge symmetry directly into the network architecture. These models integrate gauge equivariance into both convolutional layers and bilinear operations used to compute Wilson loops, which are gauge-invariant observables from lattice gauge theory. This framework was subsequently extended in [19] to include global lattice symmetries, such as rotations and reflections, thereby achieving a comprehensive symmetry-aware neural network design.